# Why Machine Learning Models Systematically Underestimate Extreme Values

### v1

@@ 1 Introduction
The advent of modern astronomical surveys has revolutionized our understanding of the universe. Large-scale observational programs have amassed unprecedented volumes of high-quality data, enabling detailed studies across multiple astronomical domains (Ivezić et al., 2012; Weinberg et al., 2013; LSST Science Collaboration et al., 2017; Aerts, 2021). This wealth of information has motivated the development of data-driven models and machine learning techniques to extract meaningful insights from complex astronomical phenomena (Baron, 2019; Huertas-Company & Lanusse, 2023; Smith & Geach, 2023).
@@
Even for more routine applications, such as determining basic stellar properties, data-driven methods have proven valuable (Fabbro et al., 2018; Leung & Bovy, 2019; Anders et al., 2023). This is particularly true for low-resolution spectra, where theoretical spectral models often fall short. A common approach is to transfer stellar parameters and elemental abundances measured from high-resolution surveys like APOGEE to lower-resolution surveys such as RAVE (Casey et al., 2017; Guiglion et al., 2020), LAMOST (Ho et al., 2017; Ting et al., 2017; Xiang et al., 2019; Zhang et al., 2020; Li et al., 2022), DESI (Zhang et al., 2024), or Gaia (Angelo et al., 2024; Hattori, 2024; Khalatyan et al., 2024; Li et al., 2024). These low-resolution observations, especially of distant stars or those obtained as part of larger cosmological surveys, typically achieve lower spectral resolution and signal-to-noise ratios (SNR).
@@
In astronomical data analysis, mapping between observables (e.g., spectra) and physical parameters (e.g., stellar properties) presents no inherent distinction between independent and dependent variables. Two main approaches exist in this field. The first approach uses generative models (e.g., The Payne, The Cannon, Cycle-StarNet) that map labels to observables and use $\chi^{2}$ minimization or MCMC sampling to obtain labels (Dafonte et al., 2016; Rix et al., 2016; Ting et al., 2016; Chandra et al., 2020; Straumit et al., 2022; Wang et al., 2022b; Laroche & Speagle, 2024; Ting et al., 2019; Różański et al., 2024; Ness et al., 2015; O’Briain et al., 2021). The second approach employs discriminative models (e.g., AstroNN, StarNet) that directly map observables to labels (Bailer-Jones et al., 1997; Snider et al., 1999; Bialek et al., 2020; Garraffo et al., 2021; Ambrosch et al., 2023; Gebran et al., 2023; Zhang et al., 2023; Fallows & Sanders, 2024; Leung & Bovy, 2019; Fabbro et al., 2018).
@@
A persistent challenge in data-driven models is systematic bias in the predictions compared to ground truth. This bias typically manifests as predictions being too low at high values and too high at low values (e.g., see the right panel of Figure 1). While this bias has anecdotally been attributed to either inadequate model training or sample imbalance leading to regression to the mean, this study demonstrates that both explanations are incorrect. We show that this bias persists even with uniformly distributed training data in both feature and label spaces with perfect ground truth labels. Furthermore, the bias depends only on the input measurement uncertainties, which systematically attenuate the regression coefficients and, consequently, the inferred label.
In linear regression models, this systematic effect, known as attenuation bias or regression dilution, has been studied analytically (see Fuller, 1987; MacMahon et al., 1990; Rosner et al., 1990; Carroll et al., 1995; Frost & Thompson, 2002, for statistical theory and applications across various fields). While these seminal works established rigorous mathematical foundations, particularly in epidemiological applications, their scope was often focused on specific use cases of their time. The emergence of modern machine learning methods and high-dimensional data analysis presents new challenges and contexts that warrant revisiting these fundamental insights.
In this study, we aim to develop a more general framework, with particular emphasis on multivariate scenarios with correlated features and large feature spaces - regimes that are increasingly relevant to modern astronomical data analysis. While previous works provided thorough mathematical treatments, our framework specifically examines how feature correlations and dimensionality interact with attenuation bias, offering insights particularly pertinent to contemporary astronomical applications where high-dimensional, correlated measurements are common. To our knowledge, this phenomenon and its implications have not been systematically examined in an astronomical context. This study therefore aims to bridge this gap by providing both theoretical understanding and practical insights specifically tailored to the astronomical community, while building upon and extending these classical statistical results.
@@ 2 Attenuation Bias
Attenuation bias, also known as regression dilution or errors-in-variables bias, systematically underestimates regression coefficients when independent variables contain measurement errors (Fuller, 1987; Carroll et al., 1995). While classical regression assumes error-free inputs with all uncertainty contained in the dependent variable, real astronomical observations invariably include measurement uncertainties in both dependent and independent variables. These systematic and statistical uncertainties can arise from multiple sources, including instrumental noise from thermal effects and detector imperfections, photon statistics, calibration errors in wavelength and flux measurements, atmospheric distortions in ground-based observations (Stubbs & Tonry, 2006; Stubbs et al., 2007; Burke et al., 2010; Noll et al., 2012; Li et al., 2016), and uncertainties in derived stellar parameters used as training labels.
@@
Figure 1: Visualization of attenuation bias through density plots comparing predicted versus true values. Left panel shows predictions with $\sigma_{\text{range}}/\sigma_{x}=10$, where the fitted regression line between the predicted $y$ as a function of true $y$ (red solid, $\lambda_{y}=0.99$) closely follows the perfect prediction line (black dashed) with only 1% bias. Right panel shows the case where measurement uncertainties are comparable to signal variance ($\sigma_{\text{range}}/\sigma_{x}=1$), where the fitted line ($\lambda_{y}=0.5$) demonstrates substantial attenuation bias, manifesting as systematic underestimation at high values and overestimation at low values.
@@ 2.1 Theoretical Framework – Univariate Linear Regression
Recall that for linear regression, the ordinary least squares (OLS) estimator for $\hat{\beta}$ when regressing $y_{\text{o}bs}$ on $x_{\text{obs}}$ has an analytic solution for its expected value:
$\mathbb{E}[\hat{\beta}]=\frac{\operatorname{Cov}(x_{\text{obs}},y_{\text{o}bs})}{\operatorname{Var}(x_{\text{obs}})}=\frac{\sum_{i}(x_{\text{obs},i}-\overline{x}_{\text{obs}})(y_{{\text{o}bs},i}-\overline{y}_{\text{o}bs})}{\sum_{i}(x_{\text{obs},i}-\overline{x}_{\text{obs}})^{2}}$ (6)
@@
It’s worth noting that, although in the derivation above, we assume the simple case of OLS where we ignore the errors in the dependent variable. However, even adopting Generalized Least Squares (GLS) to account for known heteroscedasticity in the dependent variable would not alter the conclusions of this paper (see Appendix A for details).
Recall that, in machine learning applications, our goal is to train a model on a training dataset and then use it to make predictions on new, unseen data. In astronomy, this typically involves learning relationships from a training set with known physical parameters and applying the trained model to make predictions for new observations. In the context of linear regression presented above, the linear coefficient plays the role of the “machine” (like weights in a neuron in a neural network). However, since we have established that the coefficient is biased, even after successfully training a model to minimize empirical error on the training set, predictions on new data will systematically deviate from the true values for $y$.
@@
For a concrete example, consider a high-resolution spectroscopic survey with resolution $R\equiv\lambda/\Delta\lambda=24,000$, where $\lambda$ is the wavelength and $\Delta\lambda$ is the wavelength resolution. In normalized spectra, the intrinsic variation in flux typically spans at most 10% or $\sigma_{\text{range}}=0.1$ (see also in Section 3). Even with a high signal-to-noise ratio of SNR = 100 ($\sigma_{x}=0.01$), yielding $\sigma_{\text{range}}/\sigma_{x}=10$, Equation 11 predicts $\lambda_{\beta}=\lambda_{y}=0.99$, corresponding to a 1% bias. For more typical observations, especially of distant targets in surveys like DESI and LAMOST where SNR = 10 ($\sigma_{x}=0.1$), the bias becomes severe with $\lambda_{\beta}=\lambda_{y}=0.5$, indicating a 50% underestimation. Even at moderate SNR = 30, we still find $\lambda_{\beta}=\lambda_{y}=0.9$, implying a 10% bias. These calculations demonstrate that attenuation bias is non-negligible in spectroscopic machine learning applications even under optimistic conditions, motivating a deeper investigation of this effect.
@@
Figure 2: Impact of measurement uncertainties on coefficient attenuation bias in univariate regression. The attenuation factor for the regression coefficient ($\lambda_{\beta}$) as a function of $\sigma_{\text{range}}/\sigma_{x}$ shows analytical predictions (lines) and numerical simulations (points) for linear (blue-solid), quadratic (orange-dashed), and cubic (green-dotted) terms. For each polynomial order, $\lambda_{\beta}$ is defined as the attenuation bias factor of the highest order coefficient. Linear regression coefficient simulation results follow theoretical predictions perfectly. For higher-order terms, the simulations deviate from theory due to non-linear effects, motivating our focus on the linear case where exact analytic treatment is possible. Nevertheless, the key insight remains: even with small measurement uncertainties ($\sigma_{\text{range}}/\sigma_{x}=10$), the bias is present at the percent level in the linear case and becomes more substantial in higher-order terms, suggesting that attenuation bias affects more flexible machine learning methods even though analytical solutions may not be tractable.
Figure 3: Comparison of attenuation factor $\lambda_{y}$ between numerical simulations (left) and theoretical predictions (right). The horizontal axis spans $\sigma_{\text{range}}/\sigma_{x}$ from 0.1 to 10 logarithmically, representing the ratio between true scatter and measurement uncertainty in $x$. The vertical axis covers measurement uncertainties in $y$ from 0.7 to 70 logarithmically, spanning two orders of magnitude in terms of the ratio between measurement error and the range of the data, similar to the $x$-axis range. The color scale represents the attenuation factor $\lambda_{y}$, computed by regressing predicted values against true values, with darker colors indicating stronger attenuation. The vertical gradient pattern in both panels confirms that the attenuation bias depends solely on $\sigma_{\text{range}}/\sigma_{x}$, while remaining independent of the measurement uncertainty in $y$, validating the theoretical prediction that $\lambda_{y}=\lambda_{\beta}=1/(1+(\sigma_{x}/\sigma_{\text{range}})^{2})$.
2.2 Numerical Validations – Univariate Linear Regression
@@
To validate and illustrate our theoretical analysis of attenuation bias in the univariate case, we conducted numerical simulations examining its relationship with measurement uncertainties. We generated synthetic data using a true slope $\beta=2.0$ and sampled $x_{\text{true}}$ uniformly from $[-5,5]$. We then added Gaussian measurement errors $\delta_{x}\sim\mathcal{N}(0,\sigma^{2}_{x})$ to create $x_{\text{obs}}$. The choice of a uniform distribution for $x_{\text{true}}$ was deliberate to demonstrate that attenuation bias persists even with a balanced, unbiased sample, with no uncertainty in the training label $y$ and a straightforward one-dimensional relationship. This confirms that the bias stems solely from measurement uncertainties in the independent variable rather than from sample imbalance or complex variable interactions.
@@
Although our results show the bias is independent of sample size, we simulated 10,000 samples to better visualize the effect at different measurement uncertainty ratios. To generate the predictions shown in Figure 1, we first fitted a linear regression using the observed training data ($x_{\text{obs}}$, $y_{\text{obs}}$) to obtain estimated coefficients $\hat{\beta}$, then used these coefficients to compute predicted values for the validation set. By regressing these predictions against true values, we obtain $\lambda_{y}$ that quantifies the attenuation bias in the label $y$. The results demonstrate how the magnitude of measurement uncertainties affects predictions: with $\sigma_{\text{range}}/\sigma_{x}=10$, the fitted regression line shows only 1% bias ($\lambda_{y}=0.99$) compared to the perfect prediction line. When measurement uncertainties become comparable to the signal variance ($\sigma_{\text{range}}/\sigma_{x}=1$), we observe substantial attenuation ($\lambda_{y}=0.5$) where extreme values are systematically pulled toward the mean. Recall that, for linear regression, the theoretical derivation shows that the attenuation factor for the coefficient $\lambda_{\beta}$ coincides with the attenuation factor for the predictions $\lambda_{y}$.
@@
Figure 3 extends our analysis to examine the attenuation in predictions ($\lambda_{y}$) across different measurement uncertainties in both $x$ and $y$. Following the approach in Figure 1, we computed $\lambda_{y}$ by performing linear regression on the predicted values of $y$ against true values of $y$. The horizontal axis spans $\sigma_{\text{range}}/\sigma_{x}$ from 0.1 to 10 logarithmically, while the vertical axis covers measurement uncertainties in $y$ from 0.7 to 70, with both axes spanning two orders of magnitude. The true values of $y$ span from 0 to 25 with a standard deviation of approximately 7, meaning we examine scenarios where the ratio of measurement uncertainties to true spread in $y$ is comparable to $\sigma_{\text{range}}/\sigma_{x}$ in the $x$-axis. As predicted by the theoretical framework, the vertical gradient pattern in both panels demonstrates that the attenuation bias depends solely on the ratio $\sigma_{\text{range}}/\sigma_{x}$, while remaining independent of the measurement uncertainty $\sigma_{y}$ in the dependent variable.
@@ 2.3 Attenuation Bias for Higher-Order Regression
While the theoretical framework precisely describes attenuation bias in linear regression, many astronomical applications, particularly in spectral analysis, employ more complex models (Ting et al., 2019; Fabbro et al., 2018; Leung & Bovy, 2019; O’Briain et al., 2021). For example, The Cannon employs quadratic spectral features, while The Payne and AstroNN use neural networks. Empirically, attenuation bias has been observed across various machine learning applications in astronomy, from generalized linear regression with extended features to more sophisticated approaches (see for example fig 7 and 8 of Xiang et al., 2019). Different applications exhibit varying degrees of attenuation bias, raising the question of how our framework extends to these cases.
@@
Also notably, for linear regression ($n=1$), the simulated attenuation factors (blue points) match the theoretical predictions (blue solid line) exactly across all values of $\sigma_{\text{range}}/\sigma_{x}$. However, for higher-order polynomials ($n=2,3$), while the simulations follow the general trend predicted by theory, they show increasing deviation from theoretical predictions (orange dashed and green dotted lines) as measurement uncertainties become larger ($\sigma_{\text{range}}/\sigma_{x}<1$). These deviations are expected because our non-linear approximation above relies on a Taylor expansion that assumes $\sigma_{x}/\sigma_{\text{range}}$ is small, allowing us to neglect higher-order terms. When this assumption breaks down, additional terms in the expansion become significant, leading to the observed discrepancies.

### v2

@@ Abstract
keywords:
attenuation bias—regression dilution—machine learning—stellar spectroscopy—measurement uncertainties—spectroscopic surveys—stellar parameters—data-driven models—linear regression
{CJK*}
UTF8gbsn
@@ 1 Introduction
The advent of modern astronomical surveys has revolutionized our understanding of the universe. Large-scale observational programs have amassed unprecedented volumes of high-quality data, enabling detailed studies across multiple astronomical domains (Ivezić et al., 2012; Weinberg et al., 2013; LSST Science Collaboration et al., 2017; Aerts, 2021). This wealth of information has motivated the development of data-driven models and machine learning techniques to extract meaningful insights from complex astronomical phenomena (Baron, 2019; Huertas-Company and Lanusse, 2023; Smith and Geach, 2023).
@@
Even for more routine applications, such as determining basic stellar properties, data-driven methods have proven valuable (Fabbro et al., 2018; Leung and Bovy, 2019; Anders et al., 2023). This is particularly true for low-resolution spectra, where theoretical spectral models often fall short. A common approach is to transfer stellar parameters and elemental abundances measured from high-resolution surveys like APOGEE to lower-resolution surveys such as RAVE (Casey et al., 2017; Guiglion et al., 2020), LAMOST (Ho et al., 2017; Ting et al., 2017; Xiang et al., 2019; Zhang et al., 2020; Li et al., 2022), DESI (Zhang et al., 2024), or Gaia (Angelo et al., 2024; Hattori, 2024; Khalatyan et al., 2024; Li et al., 2024). These low-resolution observations, especially of distant stars or those obtained as part of larger cosmological surveys, typically achieve lower spectral resolution and signal-to-noise ratios (SNR).
@@
In astronomical data analysis, mapping between observables (e.g., spectra) and physical parameters (e.g., stellar properties) presents no inherent distinction between independent and dependent variables. Two main approaches exist in this field. The first approach uses generative models (e.g., The Payne, The Cannon, Cycle-StarNet) that map labels to observables and use $\chi^{2}$ minimization or MCMC sampling to obtain labels (Dafonte et al., 2016; Rix et al., 2016; Ting et al., 2016; Chandra et al., 2020; Straumit et al., 2022; Wang et al., 2022a; Laroche and Speagle, 2024; Ting et al., 2019; Różański et al., 2024; Ness et al., 2015; O’Briain et al., 2021). The second approach employs discriminative models (e.g., AstroNN, StarNet) that directly map observables to labels (Bailer-Jones et al., 1997; Snider et al., 1999; Bialek et al., 2020; Garraffo et al., 2021; Ambrosch et al., 2023; Gebran et al., 2023; Zhang et al., 2023; Fallows and Sanders, 2024; Leung and Bovy, 2019; Fabbro et al., 2018).
@@
A persistent challenge in data-driven models is systematic bias in the predictions compared to ground truth. This bias typically manifests as predictions being too low at high values and too high at low values (e.g., see the right panel of Figure 1). In this study, we aim to demonstrate that this bias is not due to model training or sample distribution issues, as it persists even with uniformly distributed training data in both feature and label spaces with perfect ground truth labels. We show that the bias depends only on the input measurement uncertainties, which systematically attenuate the regression coefficients and, consequently, the inferred label.
In linear regression models, this systematic effect, known as attenuation bias or regression dilution, has been studied analytically. Fuller (1987) provided a comprehensive mathematical treatment of measurement error models, establishing the fundamental relationship for univariate cases. Subsequent work by MacMahon et al. (1990) and Rosner et al. (1990) extended these concepts to epidemiological applications, while Carroll et al. (1995) developed methods for nonlinear measurement error models. Frost and Thompson (2002) provided accessible reviews for practitioners. In astronomy, Kelly (2007, 2012) and Loredo (2004) developed methods to handle measurement errors and source uncertainties in survey analysis, though these works primarily examined univariate relationships or treated multivariate cases without explicitly analyzing the role of feature correlations in mitigating bias.
The emergence of modern machine learning methods and high-dimensional data analysis presents new challenges that warrant revisiting these fundamental insights. Specifically, our work extends the classical framework in several ways: First, we examine the low signal-to-noise regime, which, while extreme for epidemiological studies, is typical in astronomical applications. Second, we demonstrate how attenuation bias scales also applies to nonlinear transformations, which is relevant for modern machine learning architectures. Third, we provide analytical treatment for multivariate cases with correlated features—a scenario largely unexplored in both classical and astronomical literature but ubiquitous in spectroscopic analysis where physical parameters generate correlated spectral features. Finally, we provide the first systematic examination of these effects in the context of astronomical machine learning, with particular focus on stellar spectroscopy, where the combination of high dimensionality (hundreds to thousands of spectral pixels), significant measurement uncertainties, and naturally correlated features creates a unique statistical regime that directly impacts modern data-driven approaches to stellar parameter estimation.
Figure 1: Visualization of attenuation bias through density plots comparing predicted versus true values. Left panel shows predictions with $\sigma_{\text{range}}/\sigma_{x}=10$, where the fitted regression line between the predicted $y$ as a function of true $y$ (red solid, $\lambda_{y}=0.99$) closely follows the perfect prediction line (black dashed) with only 1% bias. Right panel shows the case where measurement uncertainties are comparable to signal variance ($\sigma_{\text{range}}/\sigma_{x}=1$), where the fitted line ($\lambda_{y}=0.5$).
@@ 2 Attenuation Bias
Attenuation bias, also known as regression dilution or errors-in-variables bias, systematically underestimates regression coefficients when independent variables contain measurement errors (Fuller, 1987; Carroll et al., 1995). While classical regression assumes error-free inputs with all uncertainty contained in the dependent variable, real astronomical observations invariably include measurement uncertainties in both dependent and independent variables. These systematic and statistical uncertainties can arise from multiple sources, including instrumental noise from thermal effects and detector imperfections, photon statistics, calibration errors in wavelength and flux measurements, atmospheric distortions in ground-based observations (Stubbs and Tonry, 2006; Stubbs et al., 2007; Burke et al., 2010; Noll et al., 2012; Li et al., 2016), and uncertainties in derived stellar parameters used as training labels.
@@ 2.1 Theoretical Framework – Univariate Linear Regression
For simplicity, we apply the ordinary least squares (OLS) estimator when regressing $y_{\text{o}bs}$ on $x_{\text{obs}}$, though we note that using Generalized Least Squares (GLS) to properly weight both uncertainties would not materially change our conclusions (see Appendix A). The OLS estimator has an analytic solution for its expected value:
$\mathbb{E}[\hat{\beta}]=\frac{\operatorname{Cov}(x_{\text{obs}},y_{\text{o}bs})}{\operatorname{Var}(x_{\text{obs}})}$ (6)
@@
Recall that, in machine learning applications, our goal is to train a model on a training dataset and then use it to make predictions on new, unseen data. In astronomy, this typically involves learning relationships from a training set with known physical parameters and applying the trained model to make predictions for new observations. In the context of linear regression presented above, the linear coefficient plays the role of the “machine” (like weights in a neuron in a neural network).
To elaborate on this connection: just as neural networks learn by adjusting their weights to minimize prediction error, linear regression learns by finding the optimal coefficient that minimizes the squared error between predictions and true values. The coefficient, like neural network weights, is the parameter that encodes what the model has learned from the training data and determines how the model will transform inputs into predictions. However, since we have established that the coefficient is biased, even after successfully training a model to minimize empirical error on the training set, predictions on new data will systematically deviate from the true values for $y$.
@@
For a concrete example, consider a high-resolution spectroscopic survey with resolution $R\equiv\lambda/\Delta\lambda=24,000$, where $\lambda$ is the wavelength and $\Delta\lambda$ is the wavelength resolution. Here, $x$ represents the normalized flux values in stellar spectra (i.e., spectra that have been divided by their continuum level to have a mean of approximately 1). In normalized spectra, the intrinsic variation in flux typically spans at most 10% or $\sigma_{\text{range}}=0.1$ (see also in Section 3). Even with a high signal-to-noise ratio of SNR = 100 ($\sigma_{x}=0.01$), yielding $\sigma_{\text{range}}/\sigma_{x}=10$, Equation 11 predicts $\lambda_{\beta}=\lambda_{y}=0.99$, corresponding to a 1% bias. For more typical observations, especially of distant targets in surveys like DESI and LAMOST where SNR = 10 ($\sigma_{x}=0.1$), the bias becomes severe with $\lambda_{\beta}=\lambda_{y}=0.5$, indicating a 50% underestimation. Even at moderate SNR = 30, we still find $\lambda_{\beta}=\lambda_{y}=0.9$, implying a 10% bias. These calculations demonstrate that attenuation bias is non-negligible in spectroscopic machine learning applications even under optimistic conditions, motivating a deeper investigation of this effect.
@@
2.2 Numerical Demonstrations – Univariate Linear Regression
@@
To illustrate our theoretical analysis of attenuation bias in the univariate case, we conducted numerical simulations examining its relationship with measurement uncertainties. We generated synthetic data using a true slope $\beta=2.0$ and sampled $x_{\text{true}}$ uniformly from $[-5,5]$. We then added Gaussian measurement errors $\delta_{x}\sim\mathcal{N}(0,\sigma^{2}_{x})$ to create $x_{\text{obs}}$. The choice of a uniform distribution for $x_{\text{true}}$ was deliberate to demonstrate that attenuation bias persists even with a balanced, unbiased sample, with no uncertainty in the training label $y$ and a straightforward one-dimensional relationship. This confirms that the bias stems solely from measurement uncertainties in the independent variable rather than from sample imbalance or complex variable interactions.
@@
Although our results show the bias is independent of sample size, we simulated 10,000 samples to better visualize the effect at different measurement uncertainty ratios. Unless stated otherwise, we adopt an 8:2 split for training and validation throughout this study. To generate the predictions shown in Figure 1, we first fitted a linear regression using the observed training data ($x_{\text{obs}}$, $y_{\text{obs}}$) to obtain estimated coefficients $\hat{\beta}$, then used these coefficients to compute predicted values for the validation set.
By regressing these predictions against true values, we obtain $\lambda_{y}$ that quantifies the attenuation bias in the label $y$. The results demonstrate how the magnitude of measurement uncertainties affects predictions: with $\sigma_{\text{range}}/\sigma_{x}=10$, the fitted regression line shows only 1% bias ($\lambda_{y}=0.99$) compared to the perfect prediction line. When measurement uncertainties become comparable to the signal variance ($\sigma_{\text{range}}/\sigma_{x}=1$), we observe substantial attenuation ($\lambda_{y}=0.5$) where extreme values are systematically pulled toward the mean. Recall that, for linear regression, the theoretical derivation shows that the attenuation factor for the coefficient $\lambda_{\beta}$ coincides with the attenuation factor for the predictions $\lambda_{y}$.
@@
Through additional simulations varying both $x$ and $y$ measurement uncertainties, we confirmed that the attenuation bias depends solely on the ratio $\sigma_{\text{range}}/\sigma_{x}$ while remaining independent of measurement uncertainty in $y$, exactly as predicted by our theoretical framework.
@@ 2.3 Attenuation Bias for Higher-Order Regression
While the theoretical framework precisely describes attenuation bias in linear regression, many astronomical applications, particularly in spectral analysis, employ more complex models (Ting et al., 2019; Fabbro et al., 2018; Leung and Bovy, 2019; O’Briain et al., 2021). For example, The Cannon employs quadratic spectral features, while The Payne and AstroNN use neural networks. Empirically, attenuation bias has been observed across various machine learning applications in astronomy, from generalized linear regression with extended features to more sophisticated approaches (see for example fig 7 and 8 of Xiang et al., 2019). Different applications exhibit varying degrees of attenuation bias, raising the question of how our framework extends to these cases.
@@
Figure 2: Impact of measurement uncertainties on coefficient attenuation bias in univariate regression. The attenuation factor for the regression coefficient ($\lambda_{\beta}$) as a function of $\sigma_{\text{range}}/\sigma_{x}$ shows analytical predictions (lines) and numerical simulations (points) for linear (blue-solid), quadratic (orange-dashed), and cubic (green-dotted) terms. For each polynomial order, $\lambda_{\beta}$ is defined as the attenuation bias factor of the highest order coefficient.
@@
Also notably, for linear regression ($n=1$), the simulated attenuation factors (blue points) match the theoretical predictions (blue solid line) exactly across all values of $\sigma_{\text{range}}/\sigma_{x}$. However, for higher-order polynomials ($n=2,3$), while the simulations follow the general trend predicted by theory, they show increasing deviation from theoretical predictions (orange dashed and green dotted lines) as measurement uncertainties become larger ($\sigma_{\text{range}}/\sigma_{x}<1$). The simulations reveal that the actual attenuation effects are even stronger than predicted by our simplified theoretical model in these regimes. These deviations are expected because our non-linear approximation above relies on a Taylor expansion that assumes $\sigma_{x}/\sigma_{\text{range}}$ is small, allowing us to neglect higher-order terms. When this assumption breaks down, additional terms in the expansion become significant, leading to the observed discrepancies.